An ideal gas that occupies 1.2m3 at a pressure of 1 x 105 Pa and a temperature of 27C is compressed to a volume of 0.6m3 and heated to a temperature of 227C. What is the new pressure? 


1
Expert's answer
2022-04-28T16:42:19-0400
"\frac{P_1V_1}{T_1}=\frac{P_2V_2}{T_2}"

"P_2=P_1\frac{V_1T_2}{V_2T_1}"

"P_2=1.0*10^5\times\frac{1.2*(27+273)}{0.6*(227+273)}=1.2*10^5\:\rm Pa"